OpenStudy (anonymous):

conditions for using the normal approximation for a sampling distribution of means?

OpenStudy (anonymous):

@ash2326 @hartnn @ZeHanz @TuringTest @sauravshakya anyone know?

OpenStudy (ash2326):

If the no. of means approaches a large number, then the distribution can be approximated as Normal Distribution.

OpenStudy (anonymous):

ok thats what i had ! thanks! do u know this one?

OpenStudy (anonymous):

Describe the sampling distribution of x-bar for samples of size 20 drawn from a normal population with a mean of 50 and standard deviation of 4.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

We need to describe it, maybe we have to find it's probability distribution function pdf

OpenStudy (ash2326):

It has to be normal obviously mean we have 50 and standard deviation 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what else?

OpenStudy (ash2326):

\[\large f(x)=\frac 1 {\sqrt {2 \pi} \sigma}\times e^{\frac{x-\mu}{2\sigma^2} }\] put \[\sigma =4\] \[\mu =50\]

OpenStudy (ash2326):

@schmidtdancer I think probably I'm correct. But do get somebody to check it once

OpenStudy (anonymous):

So what dhould i put as my answer?

OpenStudy (ash2326):

f(x)

OpenStudy (ash2326):

for \[-\infty<x<\infty\]

OpenStudy (anonymous):

sorry lol, what do i put?

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

im confused on what to put for my answer

OpenStudy (ash2326):

Put the distribution saying the distribution is approximated to normal distribution

OpenStudy (anonymous):

all i have is It has to be normal. The mean is 50 and the standard deviation is 4.

OpenStudy (ash2326):

yes, put the distribution f(x) what I have posted

OpenStudy (anonymous):

Can u repost?

OpenStudy (ash2326):

\[\large f(x)=\frac 1 {\sqrt {2 \pi} \sigma}\times e^{\frac{x-\mu}{2\sigma^2} }\] \[\sigma=4\] \[\mu=50\]

OpenStudy (anonymous):

how do i write that in word form? @ash2326

OpenStudy (anonymous):

I got this f(x) = 1 / sqrt(2pi*mean) i cant figure out the rest

OpenStudy (ash2326):

* exponential( (x^2-mean)/(2*standard deviation^2))

OpenStudy (anonymous):

This ? f(x) = 1 / sqrt(2pi*mean)* exponential(x^2-mean)/(2*standard deviation^2))

OpenStudy (anonymous):

@ash2326 ?

OpenStudy (ash2326):

Yes, sorry my net was giving me problems